Factorise the following expressions
-4a^2 +4ab–4 ca

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

-4a^2 +4ab–4 ca
- 4a^2= -1×2×2×a×a
4ab = 2×2×a×b
- 4ca = -1×2×2×c×a
Common factor of - 4a^2, 4ab , - 4ca are 2, 2, a i.e. 4a
So,
- 4a^2+4 ab-4 ca = 4a(-a+b-c)

Thanks